(x^2+y^2)p+2xyq=dz/(x+y)z


1
Expert's answer
2021-03-07T16:45:05-0500

The Langrage equation is:


"(x^2+y^2)p+2xyq=(x+y)z"

From the general equaton:


"Pp + Qq = R"

From the above equation,

"P = x^2+ y^2\\
Q = 2xy\\
R = (x+y)z"


From the auxilliary equation,


"\dfrac{dx}{P} = \dfrac{dy}{Q} = \dfrac{dz}{R}"

Hence:


"\dfrac{dx}{x^2+y^2} = \dfrac{dy}{2xy} = \dfrac{dz}{(x+y)z} \qquad \cdots \cdots \cdots eqn(i)\\"

From the first two and last ratios and solve, we have:


"\dfrac{dx+dy}{(x^2+y^2)+2xy}=\dfrac{dz}{(x+y)z}\\"


Recall:


"x^2+y^2 = (x+y)^2 - 2xy \qquad \cdots eqn (*) \\""\therefore \dfrac{dx+dy}{(x+y)^2 -2xy +2xy}=\dfrac{\frac{dz}{z}}{x+y}\\
\dfrac{dx+dy}{(x+y)^2}=\dfrac{\frac{dz}{z}}{x+y}\\
\dfrac{dx+dy}{x+y}=\dfrac{dz}{z}\\
\dfrac{d(x+y)}{x+y}=\dfrac{dz}{z}\\"


Integrating both sides:


"\int \dfrac{d(x+y)}{x+y}=\int \dfrac{dz}{z}\\
\ln (x+y) = \ln z +\ln c\\
\ln (x+y) =\ln cz\\
x+y = cz\\
\implies c_1 = \dfrac{x+y}{z} \qquad \cdots\cdots eqn (ii)"

Consider the last ratio in eqn (i):


"\frac{dz}{(x+y)z}"

From eqn (ii)


"z = \dfrac{x+y}{c_1}"

Substituting the value of "z" in the last ratio:


"\frac{dz}{(x+y)\frac{(x+y)}{c_1}}\\
\frac{dz}{\frac{(x+y)^2}{c_1}}\\
\implies \frac{c_1dz}{(x+y)^2} \qquad \cdots eqn (iii)"

By replacing the third ration in eqn (i) with eqn(iii)


"\frac{dx}{x^2+y^2} = \frac{dy}{2xy} = \frac{c_1dz}{(x+y)^2}"

Thus:


"\frac{dx-dy-c_1dz}{x^2+y^2-2xy- (x+y)^2} = k \qquad \text{(say)}"

Using eqn (*):


"\frac{dx-dy-c_1dz}{x^2+y^2-2xy- [x^2+y^2-2xy]} = k\\
\frac{dx-dy-c_1dz}{x^2+y^2-2xy- x^2-y^2+2xy} = k\\
\frac{dx-dy-c_1dz}{0} = k\\
\implies dx-dy-c_1dz = 0\\"

Integrating through:


"\int dx-\int dy-\int c_1dz = \int 0\\
x-y-c_1z=c_2\\"

Recall:

"c_1 = \dfrac{x+y}{z}"

"\therefore x-y-\Big(\frac{x+y}{z}\Big)z=c_2\\
x-y-x-y=c_2\\
\implies c_2 = -2y"

Therefore, the solution is:


"f(u,v) = \Big(\dfrac{x+y}{z},\; -2y\Big) = 0"